40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 27 23:49:56 2014, Koji, Update, General, IR beam found, TTs not aligned well yet 
    Reply  Mon Jun 30 16:55:24 2014, ericq, Update, General, IFO (mostly) aligned 
       Reply  Mon Jun 30 22:06:55 2014, Jenne, Update, ASC, IFO (mostly) aligned - ASS working 
          Reply  Wed Jul 2 16:18:32 2014, Jenne, Update, ASC, IFO aligned, PRMI + 2 arms achieved 
Message ID: 10114     Entry time: Mon Jun 30 22:06:55 2014     In reply to: 10113     Reply to this: 10124
Author: Jenne 
Type: Update 
Category: ASC 
Subject: IFO (mostly) aligned - ASS working 

[Koji, Jenne]

The Yarm ASS is now working (as is the Xarm ASS).  Both of the TT's pitch servos had a sign flip.  We don't know why.

To start, we lowered the matrix elements that push on the TTs by a factor of 3, to compensate for the new factor of 3 in the slider gains:  ezcastep C1:ASS-YARM_OUT_MTRX_5_5 /3 C1:ASS-YARM_OUT_MTRX_5_7 /3 C1:ASS-YARM_OUT_MTRX_6_6 /3 C1:ASS-YARM_OUT_MTRX_6_8 /3 C1:ASS-YARM_OUT_MTRX_7_5 /3 C1:ASS-YARM_OUT_MTRX_7_7 /3 C1:ASS-YARM_OUT_MTRX_8_6 /3 C1:ASS-YARM_OUT_MTRX_8_8 /3

We turned off all 4 tip tilt ASS servos (in the Yarm ASS servo screen), and turned them on one at a time.  By doing this, we discovered that the pitch servos for both TT1 and TT2 needed to have the opposite sign from what they used to have.  However, the yaw servos kept the original signs.  It really doesn't make sense to me why this should be, but this is the way the ASS servo works.  We left both Xarm and Yarm ASSs on for several minutes, and saw that they didn't push any mirrors out of alignment.

The ASS_DOTHER_ON burt snapshot has been resaved with the new values.

 

Also, earlier this evening, I aligned the Yarm green beam to the cavity, although the cavity was not optimally aligned, so this needs to be re-done.

 

On our to-do list should be to add the tip tilt slider values to the DAQ channels list.

ELOG V3.1.3-